The table below represents the values of x (fork-lift driving experiences in years) and y (monthly insurance in australian dollar).

x

y

5

64

2

87

12

50

9

71

15

44

6

56

25

42

16

60

  1. Find the least square regression line.
  2. Calculate and critically comment on the value of r and r2.
Answer #1

(a)

Sample x y x2 y2 x.y
1 5 64 25 4,096 320
2 2 87 4 7,569 174
3 12 50 144 2,500 600
4 9 71 81 5,041 639
5 15 44 225 1,936 660
6 6 56 36 3,136 336
7 25 42 625 1,764 1,050
8 16 60 256 3,600 960
Totals 90 474 1,396 29,642 4,739
Σx Σy Σx2 Σy2 Σx.y
Sample size (n) = 8
= Σx/n = 11.3
ȳ = Σy/n = 59.3
α = Σx.y - n.x̅ .ȳ = -593.5
β = Σx2 - n. 2 = 383.5
Slope (b) = α/β = -1.548
Intercept (a) = ȳ - b. = 76.660

The regression equation will be written as ŷ = 76.66 - 1.548 * x

(b)

Sample x y x2 y2 x.y
1 5 64 25 4,096 320
2 2 87 4 7,569 174
3 12 50 144 2,500 600
4 9 71 81 5,041 639
5 15 44 225 1,936 660
6 6 56 36 3,136 336
7 25 42 625 1,764 1,050
8 16 60 256 3,600 960
Totals 90 474 1,396 29,642 4,739
Σx Σy Σx2 Σy2 Σx.y
Sample size (n) = 8
Numerator = nxy - Σxy = -4748.00
Denominator = [nx2 - (Σx)2]1/2.[ny2 - (Σy)2]1/2 = 6182.82
Correlation coefficient (r) = Numerator / Denominator -0.7679

So,

The correlation coefficient, r = -0.768

The correlation coefficient suggests that there is a strong (since > 0.5) and indirect (since negative) correlation between 'x' and 'y'.

The coefficient of determination, r2 = (-0.7679)^2 = 0.5897

This value of r2 says that about 59% of the total variation of 'y' can be explained by the independent variable 'x'. This is an indication that the explanatory power of the model is satisfactory.